Which is most difficult to monetie? a)maintenance cost, b)teling price e, prestige d.) interest on debt .) cost 2.)$100 isdeposited in a savigs account e 6% interest and an the inerest sleft what is the account balance after 3y a)5840 1000 1180 1191 e3000 act 3) You are asked for a loan and promned repayment of 5250 two years from now Ithe mmm interest rate you wศ accept 8%, what is the maimun amount you will loan? 4) The annual amount of a series of payments to be made at the end of each of the mest 12 ys. is $500 what is the present worth of he payments at 8% imerest compourold annualy? 5) Consider a prospective investment in a project having a first cost off $300,000, operating and -tenance costs of S3500yr, .nd an estimated net disposal value of SS0αο at tbe end of th yrs. Interest rate is 8%, what is the present equivalent cost of investment d pl ho-is thirty yrs? )5670,000 b)$683,000 )$720,000 5791,000 e15950,000
Answer #1

1.

C

It is the prestige that is intangible in nature. So, it is very difficult to get monetized.

2.

D

Account balance at the end of 3 years = 1000*(1+6%)^3 = $1191 approx.

3.

C

Present value of the loan = 150/(1.08^2) = $128.6 or $129

4.

B

Present value of the deposits = 500*(1-1/1.08^12)/.08 = $3768

Pl repost other unanswered questions with clear numbers mentioned so that it can be answered. You can post it in comments also so that I can answer.